# Sharp thresholds for spanning regular subgraphs

### v2

@@ Acknowledgements
By [17, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{(d)}_{n}=\mathcal{F}^{(d)}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$.
@@
Let $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and let $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{d})$ be a uniformly random $m$-subset. Then whp $(F,\mathbf{W})$ is not $\ell_{0}$-bad.
@@
Let $\delta(n)$ be a slowly decreasing function and let $t\in\{0,\ldots,2n-5d\}$. Due to Claim 6.2, we have
$\mathbb{P}((F,\mathbf{W})\text{ is $\ell_{0}$-bad},\,|\mathcal{M}(F,\mathbf{W})|<\delta(n)M(t)\mid|F\cap\mathbf{W}|=t)=o(1).$
@@
$\mathbb{P}((F,\mathbf{W})\text{ is $\ell_{0}$-bad},\,|\mathcal{M}(F,\mathbf{W})|\geq\delta(n)M(t)\mid|F\cap\mathbf{W}|=t)=o(1).$

### v3

@@ Acknowledgements
By [18, Corollary 1.16], whp there exists a multiset $\mathcal{F}^{(d)}_{n}=\mathcal{F}^{(d)}_{n}(\mathbf{G})$ of graphs of size $\lfloor\ln n\rfloor$ comprising a single subgraph $H$ of almost every $F\in\mathcal{F}_{n}$ so that $H\cup\mathbf{G}$ contains some $F^{\prime}\in\mathcal{F}_{n}$.
@@
Let $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and let $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{d})$ be a uniformly random $m$-subset. Then $(F,\mathbf{W})$ is not $\ell_{0}$-bad with probability at least $1-\exp(-\Omega(\sqrt{n}))$.
@@
Fix $\varepsilon^{\prime}>0$ small enough and let $\delta(n)=\exp(-\varepsilon^{\prime}\sqrt{n})$. Fix $t\in\{0,\ldots,2n-5d\}$. Due to Claim 6.2, we have
$\mathbb{P}((F,\mathbf{W})\text{ is $\ell_{0}$-bad},\,|\mathcal{M}(F,\mathbf{W})|<\delta(n)M(t)\mid|F\cap\mathbf{W}|=t)=\exp(-\Omega(\sqrt{n})).$
@@
$\mathbb{P}((F,\mathbf{W})\text{ is $\ell_{0}$-bad},\,|\mathcal{M}(F,\mathbf{W})|\geq\delta(n)M(t)\mid|F\cap\mathbf{W}|=t)=\exp(-\Omega(\sqrt{n})).$